OpenStudy (anonymous):

Which of the following is the best definition of viscosity? A. the tendency of a fluid to support a floating object B. the ability of a fluid to transmitted a force C. the thickness of a fluid, or its resistance to flow

OpenStudy (anonymous):

C. Is the right answer, viscosity is how a liquid resist flow. It has units of \[\frac{ N*S }{ m^2 }\] for the coefficient viscosity.

OpenStudy (anonymous):

One addition: the units are newton-seconds per meter.

OpenStudy (shamim):

The above unit is wrong!!!

OpenStudy (shamim):

U know the equation F=eta*A*dv/dy Here eta is viscocity

OpenStudy (shamim):

Can u tell me the unit of eta frm my equation

OpenStudy (shamim):

Eta=F*dy/A.dv

OpenStudy (shamim):

Unit of viscocity=N.m/(m^2.m.s^-1)=N.s/m^2